# Asymptotic theory of in-context learning by linear attention

### v1

@@ Acknowledgements
Note that the test distribution $\mathcal{P}_{\mathrm{test}}$ crucially depends on the probability distribution of the task vector $w$ used in the linear model in (1). For the ICL task, we have $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$, the uniform distribution on the sphere $\mathcal{S}^{d-1}(\sqrt{d})$. For the IDG task, $w$ is sampled uniformly from the set $\left\{w_{1},w_{2},\ldots,w_{k}\right\}$, where these $k$ vectors are the same as those used in the training data [see (7)]. In what follows, we slightly abuse the notation by writing $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$.
@@
Using the equivalent representations in (A.1) and (A.2), it is straightforward to verify the estimates of the first (conditional) moments in (SI.2.2). To show (SI.2.3), we note that
@@
Using the representation in (A.2), we have
@@
We use $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$. Recall our discussions in Section 2.6. For the ICL task, $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. It is straightforward to check that, in this case,
@@
$G(\pi)=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\pi\Omega+\tau\lambda I\right)^{-1},$ (SI.3.4)
where $\Omega\in\mathbb{R}^{(d^{2}+d)\times(d^{2}+d)}$ is a symmetric positive-semidefinite matrix and $\pi$ is a nonnegative scalar. The original resolvent $G$ in (SI.3.3) is a special case, corresponding to $\pi=0$.
@@
$\Omega_{e}=\begin{bmatrix}0&\\ &\Omega\end{bmatrix},$ (SI.3.6)
where $\Omega$ is the $(d^{2}+d)\times(d^{2}+d)$ matrix in (SI.3.4). Define an extended resolvent matrix
$G_{e}(\pi)=\frac{1}{\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Omega_{e}+\tau\lambda I}.$ (SI.3.7)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}G(\pi)=-G(\pi)\Omega G(\pi),$ (SI.3.15)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)$ $=\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\Omega G(\pi)\operatorname{vec}(H_{\nu})$ (SI.3.16)
$=q^{\top}(\pi)\Omega q(\pi).$ (SI.3.17)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr{|}_{\pi=0}$ $=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\Omega\operatorname{vec}(\Gamma^{\ast})$ (SI.3.18)
@@
$\Omega=I_{d}\otimes B_{\mathrm{test}},$ (SI.3.20)
@@
Consider the extended resolvent $G_{e}(\pi)$ in (SI.3.7), with $\Omega_{e}$ chosen in the forms of (SI.3.6) and (SI.3.20). Let $\widetilde{G}_{e}$ be another matrix of the same size as $G_{e}(\pi)$. We say that $\widetilde{G}_{e}$ and $G_{e}(\pi)$ are asymptotically equivalent, if the following conditions hold.
@@
$\mathcal{G}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\chi_{\pi}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Omega_{e}+\tau\lambda I\right)^{-1},$ (SI.4.8)
in the sense of Definition 1. In the above expression, $\Omega_{e}$ is the matrix in (SI.3.6) with $\Omega=I_{d}\otimes B_{\mathrm{test}}$.
@@
$G_{e}\left(\textstyle\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Omega_{e}+\tau\lambda I\right)=I.$ (SI.4.10)
@@
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}+G_{e}(\pi\Omega_{e}+\tau\lambda I)=I+\Delta_{1},$ (SI.4.15)
@@
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]+G_{e}(\pi\Omega_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2},$ (SI.4.17)
@@
$\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+G_{e}(\pi\Omega_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2}+\Delta_{3},$ (SI.4.21)
@@
$G_{e}\left(\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi(w_{\mu})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\pi\Omega_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.23)
@@
$G_{e}\left(\frac{\tau}{k}\sum_{i\in[k]}\frac{1}{1+\chi(w_{i})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{i}^{\top}\otimes\begin{bmatrix}w_{i}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{i}\otimes\begin{bmatrix}w_{i}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{i})\end{bmatrix}+\pi\Omega_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.24)
@@
$G_{e}\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Omega_{e}+\tau\lambda I\right)$ (SI.4.30)
@@
$\widehat{\mathcal{G}}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Omega_{e}+\tau\lambda I\right)^{-1}.$ (SI.4.31)
@@
To compute the first term on the right-hand side of the above estimate, we directly invert the block matrix $\widehat{\mathcal{G}}_{e}(\pi)$ in (SI.4.31). Recall that $\Omega_{e}$ is chosen in the forms of (SI.3.6) and (SI.3.13). It is then straightforward to verify that
@@
We consider two different types of test data distributions $\mathcal{P}_{\mathrm{test}}$: ICL and IDG. [See Section 2.6 in the main text for details.] From (SI.2.16) and (SI.2.17), these two settings correspond to choosing
@@
and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (B.3). Moreover, the scalar $c_{e}$ in (SI.5.4) is defined as
@@
In the above display, (SI.5.13) follows from (SI.3.12) and the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. The equality in (SI.5.14) is due to (SI.5.9) and (SI.5.11). To reach (SI.5.15), we note that $c(0)=e_{1}^{\top}G_{e}(0)e_{1}$ and $c^{\ast}(0)=e_{1}^{\top}\mathcal{G}_{e}(0)e_{1}$. Thus, $c(0)\simeq c^{\ast}(0)$ due to the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. In Appendix B, we show that
@@
where $F_{R}(\cdot)$ is the function defined in (B.1), and $\nu$ is the parameter given in (SI.5.8). The second equality (SI.5.21) is obtained from the explicit formula for $(E_{\mathrm{tr}}+\xi I)^{-1}$ in (B.7).
@@
By using the asymptotic characterizations given in (B.19) and (B.20), we then have
@@
Substituting (SI.5.24), (SI.5.25), and (B.21) into (SI.5.18) yields
@@
where $\nu$ is the constant defined in (SI.5.8). To reach the last step, we have used the estimate given in (B.19).
@@
where the second step follows from (B.19) and (B.20). From (B.7),
@@
where $\mathcal{M}_{\kappa}(x)$ is the function defined in (B.3). Then
@@
By combining (SI.5.41) and the characterization in (B.4), we can directly solve for $\mu$ and get $\mu^{\ast}=q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast})$ as given in (SI.5.39). The characterization in (SI.5.40) (for the case of $\tau<1$) then directly follows from (SI.5.33), (SI.5.38), and (4) after some lengthy calculations.
@@
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ $\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ (SI.5.51)
=$=\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\qquad\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\qquad\tau>1\end{cases},$ (SI.5.52)

### v2

@@ Acknowledgments
Note that the test distribution $\mathcal{P}_{\mathrm{test}}$ crucially depends on the probability distribution of the task vector $w$ used in the linear model $y=w^{\top}x+\eta$. For the ICL task, we have $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$, the uniform distribution on the sphere $\mathcal{S}^{d-1}(\sqrt{d})$. For the IDG task, $w$ is sampled uniformly from the set $\left\{w_{1},w_{2},\ldots,w_{k}\right\}$, where these $k$ vectors are the same as those used in the training data. In what follows, we slightly abuse the notation by writing $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$.
You will note that the distribution $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$ chosen above does not match the distribution $w\sim\mathcal{N}(0,I_{d}).$ For analytical tractibility and convenience, the remaining calculations will have $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$ not $w\sim\mathcal{N}(0,I_{d}).$ For large $d$ (in practice, this seems to be any $d>80$), simulations show that these choices are indistinguishable.
@@
Using the equivalent representations in (SI.7.1) and (SI.7.2), it is straightforward to verify the estimates of the first (conditional) moments in (SI.2.2). To show (SI.2.3), we note that
@@
Using the representation in (SI.7.2), we have
@@
We use $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$. Recall our discussions in Section II.6. For the ICL task, $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. It is straightforward to check that, in this case,
@@
$G(\pi)=\left(\textstyle\sum_{\mu\in[n]}\operatorname{vec}({H_{\mu}})\operatorname{vec}({H_{\mu}})^{\top}/d+\pi\Pi+\tau\lambda I\right)^{-1},$ (SI.3.4)
where $\Pi\in\mathbb{R}^{(d^{2}+d)\times(d^{2}+d)}$ is a symmetric positive-semidefinite matrix and $\pi$ is a nonnegative scalar. The original resolvent $G$ in (SI.3.3) is a special case, corresponding to $\pi=0$.
@@
$\Pi_{e}=\begin{bmatrix}0&\\ &\Pi\end{bmatrix},$ (SI.3.6)
where $\Pi$ is the $(d^{2}+d)\times(d^{2}+d)$ matrix in (SI.3.4). Define an extended resolvent matrix
$G_{e}(\pi)=\frac{1}{\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I}.$ (SI.3.7)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}G(\pi)=-G(\pi)\Pi G(\pi),$ (SI.3.15)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)$ $=\frac{1}{d^{3}}\sum_{\mu,\nu\in[n]}y_{\mu}y_{\nu}\operatorname{vec}(H_{\mu})^{\top}G(\pi)\Pi G(\pi)\operatorname{vec}(H_{\nu})$ (SI.3.16)
$=q^{\top}(\pi)\Pi q(\pi).$ (SI.3.17)
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr{|}_{\pi=0}$ $=\frac{1}{d}\left(\operatorname{vec}(\Gamma^{\ast})\right)^{\top}\Pi\operatorname{vec}(\Gamma^{\ast})$ (SI.3.18)
@@
$\Pi=I_{d}\otimes B_{\mathrm{test}},$ (SI.3.20)
@@
Consider the extended resolvent $G_{e}(\pi)$ in (SI.3.7), with $\Pi_{e}$ chosen in the forms of (SI.3.6) and (SI.3.20). Let $\widetilde{G}_{e}$ be another matrix of the same size as $G_{e}(\pi)$. We say that $\widetilde{G}_{e}$ and $G_{e}(\pi)$ are asymptotically equivalent, if the following conditions hold.
@@
$\mathcal{G}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\chi_{\pi}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1},$ (SI.4.8)
in the sense of Definition 1. In the above expression, $\Pi_{e}$ is the matrix in (SI.3.6) with $\Pi=I_{d}\otimes B_{\mathrm{test}}$.
@@
$G_{e}\left(\textstyle\sum_{\mu\in[n]}z_{\mu}z_{\mu}^{\top}+\pi\Pi_{e}+\tau\lambda I\right)=I.$ (SI.4.10)
@@
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}z_{\mu}z_{\mu}^{\top}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1},$ (SI.4.15)
@@
$\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\mathbb{E}_{w_{\mu}}\left[z_{\mu}z_{\mu}^{\top}\right]+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2},$ (SI.4.17)
@@
$\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi^{\mu}(w_{\mu})}G_{e}^{[\mu]}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+G_{e}(\pi\Pi_{e}+\tau\lambda I)=I+\Delta_{1}+\Delta_{2}+\Delta_{3},$ (SI.4.21)
@@
$G_{e}\left(\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi(w_{\mu})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.23)
@@
$G_{e}\left(\frac{\tau}{k}\sum_{i\in[k]}\frac{1}{1+\chi(w_{i})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{i}^{\top}\otimes\begin{bmatrix}w_{i}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{i}\otimes\begin{bmatrix}w_{i}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{i})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.24)
@@
$G_{e}\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)$ (SI.4.30)
@@
$\widehat{\mathcal{G}}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1}.$ (SI.4.31)
@@
To compute the first term on the right-hand side of the above estimate, we directly invert the block matrix $\widehat{\mathcal{G}}_{e}(\pi)$ in (SI.4.31). Recall that $\Pi_{e}$ is chosen in the forms of (SI.3.6) and (SI.3.13). It is then straightforward to verify that
@@
We consider two different types of test data distributions $\mathcal{P}_{\mathrm{test}}$: ICL and IDG. [See Section II.6 in the main text for details.] From (SI.2.16) and (SI.2.17), these two settings correspond to choosing
@@
and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (SI.8.3). For convenience, this can be recast so that $\xi$ is the positive double root of
$\kappa\tau\xi^{4}+(-\kappa\lambda\tau+\kappa\tau x+\kappa\tau-\kappa x-\tau^{2}-\kappa+\tau)\xi^{3}+(-\kappa\lambda\tau x-\kappa\lambda\tau+2\lambda\tau^{2}-\tau^{2}x-\lambda\tau+2\tau x-x)\xi^{2}+(-\lambda^{2}\tau^{2}+2\lambda\tau^{2}x-2\lambda\tau x)\xi-\lambda^{2}\tau^{2}x.$
Moreover, the scalar $c_{e}$ in (SI.5.4) is defined as
@@
In the above display, (SI.5.13) follows from (SI.3.12) and the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. The equality in (SI.5.14) is due to (SI.5.9) and (SI.5.11). To reach (SI.5.15), we note that $c(0)=e_{1}^{\top}G_{e}(0)e_{1}$ and $c^{\ast}(0)=e_{1}^{\top}\mathcal{G}_{e}(0)e_{1}$. Thus, $c(0)\simeq c^{\ast}(0)$ due to the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. In Appendix SI.8, we show that
@@
where $F_{R}(\cdot)$ is the function defined in (SI.8.1), and $\nu$ is the parameter given in (SI.5.8). The second equality (SI.5.21) is obtained from the explicit formula for $(E_{\mathrm{tr}}+\xi I)^{-1}$ in (SI.8.7).
@@
By using the asymptotic characterizations given in (SI.8.19) and (SI.8.20), we then have
@@
Substituting (SI.5.24), (SI.5.25), and (SI.8.21) into (SI.5.18) yields
@@
where $\nu$ is the constant defined in (SI.5.8). To reach the last step, we have used the estimate given in (SI.8.19).
@@
where the second step follows from (SI.8.19) and (SI.8.20). From (SI.8.7),
@@
where $\mathcal{M}_{\kappa}(x)$ is the function defined in (SI.8.3). Then
@@
By combining (SI.5.41) and the characterization in (SI.8.4), we can directly solve for $\mu$ and get $\mu^{\ast}=q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast})$ as given in (SI.5.39). The characterization in (SI.5.40) (for the case of $\tau<1$) then directly follows from (SI.5.33), (SI.5.38), and (4) after some lengthy calculations.
@@
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ $\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\qquad\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\qquad\tau>1\end{cases},$ (SI.5.51)